A 60 degree prism of refractive index is 1.5. what is the angle minimum deviation? calculate the angle of refraction of light passing through the prism at minimum deviation and the angle of incidence on the prism at minimum deviation
1
Expert's answer
2017-11-24T14:17:07-0500
The angle minimum deviation D can be found from the equation:
n=sin⁡〖(A+D)/2〗/sin⁡〖A/2〗
sin⁡〖(A+D)/2〗=n sin⁡〖A/2〗=1.5 sin⁡30
sin⁡〖(A+D)/2〗=n sin⁡〖A/2〗=1.5 1/2=3/4
(A+D)/2=sin^(-1)⁡〖3/4〗=48.59°
D=2(48.59)-60=37°.
The refracting angle is
r=A/2=30°.
The angle of incidence is
i=(A+D)/2=(60+37)/2=48.5°